Calculate the enthalpy of vaporization of C4H10. This compound has vapor pressures of 492 mmHg and 355 mmHg at -12.0 °C and -20.0 °C, respectively. (R = 8.314 J/mol·K)
Answer #1

Using, ln(P1/P2) = ∆Hvap/R [1/T2 - 1/T1]

T1 = -12°C = -12+273.15 K = 261.15 K

T2 = -20°C = -20+273.15 = 253.15 K

Substituting values, we get :

ln(492/355) = ∆Hvap/8.314 [1/253.15 -1/261.15]

8.314×ln(1.3859) = ∆Hvap [0.00395 - 0.003829]

2.713 = ∆Hvap×0.000121

∆Hvap = 2.713/0.000121

= 22.42 KJ/mol